The ratio of the energy of a photon of 200 Å wavelength radiation to that of 400 Å radiation is
a. 1/4
b. 4
c. 1/2
d. 2

Grade:11

1 Answers

Aditi Chauhan
9 years ago
E = hc/λ ⇒ E_1/E_2 = λ_2/λ_1 =2